# Learning to Optimize for Mixed-Integer Non-linear Programming

### v3

@@ 6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section D.2. Due to the structure of the constraints, it is straightforward to always construct a trivial feasible integer solution, ensuring that the problem remains feasible across instances.
@@
We conducted experiments on mixed-integer Rosenbrock problems with the number of decision variables ranging from 2 to 20,000; the number of constraints was fixed at 5. The results in Table 4 show that RC and LT exhibit strong performance, even outperforming EX in smaller cases, while RS performs slightly worse than RC and LT in terms of objective value. However, as the problem size increases to 10,000 variables, a noticeable decline in feasibility is observed for both RC and LT, while solver-based methods such as EX, N1, and RR fail to produce any solutions. As seen in previous experiments, RR and RL, which rely on rounding relaxations, continue to suffer from significant infeasibility issues.
@@
Table 4: Results for the Mixed-Integer Rosenbrock Problem. The number of decision variables varies from 2 to 20,000, while the number of constraints is 5. See the caption of Table 2 for details.
Method Metric 2×5 20×5 200×5 2000×5 20000×5
RC Obj Mean $23.30$ $56.99$ $505.47$ $6245.15$ $63394.52$
Obj Median $21.22$ $50.96$ $439.79$ $6500.15$ $60144.23$
% Infeasible $2\%$ $1\%$ $1\%$ $2\%$ $45\%$
Time (Sec) $0.0031$ $0.0033$ $0.0035$ $0.0042$ $0.0856$
LT Obj Mean $23.02$ $55.06$ $608.47$ $6224.79$ $66890.89$
Obj Median $21.27$ $54.96$ $600.25$ $6097.23$ $48003.27$
% Infeasible $4\%$ $0\%$ $0\%$ $3\%$ $33\%$
Time (Sec) $0.0031$ $0.0032$ $0.0035$ $0.0047$ $0.0843$
EX Obj Mean $18.11$ $63.58$ $59.0e$ 5 $101.0e$ 10
Obj Median $18.34$ $56.84$ $28.9e$ 2 $9235.17$
% Infeasible $0\%$ $0\%$ $0\%$ $3\%$
Time (Sec) $0.4512$ $60.1385$ $59.5281$ $60.1545$
RR Obj Mean $21.52$ $52.03$ $53.3e$ 5 $68.6e$ 6
Obj Median $22.36$ $50.92$ $500.55$ $7189.05$
% Infeasible $37\%$ $42\%$ $36\%$ $55\%$
Time (Sec) $0.6756$ $1.0860$ $3.1870$ $20.8632$
N1 Obj Mean $43.86$ $88.29$ $083.0e$ 08 $132.1e$ 13
Obj Median $28.37$ $78.59$ $961.20$ $9235.17$
% Infeasible $0\%$ $0\%$ $0\%$ $3\%$
Time (Sec) $0.0752$ $0.1647$ $0.3949$ $8.7531$
RL Obj Mean $58.36$ $64.23$ $616.06$ $6186.16$ $58414.16$
Obj Median $58.02$ $62.80$ $616.91$ $5787.59$ $54241.67$
% Infeasible $86\%$ $45\%$ $55\%$ $17\%$ $26\%$
Time (Sec) $0.0006$ $0.0010$ $0.0009$ $0.0010$ $0.0007$
RS Obj Mean $25.07$ $65.38$ $634.95$ $6566.85$ $72904.02$
Obj Median $25.28$ $63.86$ $619.46$ $6441.98$ $72533.57$
% Infeasible $0\%$ $2\%$ $0\%$ $0\%$ $35\%$
Time (Sec) $0.0018$ $0.0016$ $0.0017$ $0.0027$ $0.0066$
6.4 Effect of Increasing Penalty Weight
In this section, we evaluate the effect of progressively increasing penalty weight on the above benchmark problems. This experiment is inspired by penalty methods commonly used in optimization, such as the Augmented Lagrangian Method (Hestenes, 1969). The weight of the quadratic penalty term is typically increased throughout the iterations. This gradual increase in penalty helps avoid numerical difficulties caused by overly strict constraints at the early stages while enhancing constraint satisfaction as the solution converges.
Table 5: Comparison of Results for RC and LT with and without Increasing Penalty Weight on Benchmark Problems. See the caption of Table 2 for details.
Problem Size Method Original Increasing Penalty
Obj Mean Obj Median % Infeasible Obj Mean Obj Median % Infeasible
Convex Quadratic 100×100 RC $-13.589$ $-13.587$ $4\%$ $-10.660$ $-10.697$ $1\%$
LT $-13.745$ $-13.747$ $7\%$ $-11.548$ $-11.617$ $0\%$
200×200 RC $-31.206$ $-31.247$ $5\%$ $-25.350$ $-25.412$ $0\%$
LT $-31.261$ $-31.320$ $1\%$ $-25.110$ $-25.297$ $0\%$
500×500 RC $-73.570$ $-73.761$ $8\%$ $-60.040$ $-60.405$ $4\%$
LT $-73.457$ $-73.414$ $8\%$ $-57.507$ $-57.590$ $1\%$
Simple Non-Convex 100×100 RC $0.974$ $0.928$ $1\%$ $1.773$ $1.718$ $0\%$
[…]
200×200 RC $1.688$ $1.688$ $0\%$ $1.662$ $1.584$ $0\%$
LT $0.514$ $0.464$ $1\%$ $2.163$ $2.060$ $0\%$
[…]
Multi-Dim Rosenbrock 200×5 RC $505.47$ $439.79$ $1\%$ $519.85$ $467.59$ $0\%$
LT $608.47$ $600.25$ $0\%$ $618.51$ $624.10$ $0\%$
2000×5 RC $6245.15$ $6500.15$ $2\%$ $7276.96$ $7460.26$ $0\%$
LT $6224.79$ $6097.23$ $3\%$ $5780.85$ $5590.29$ $0\%$
20000×5 RC $63394.52$ $60144.23$ $45\%$ $63626.35$ $58977.61$ $29\%$
LT $66890.89$ $48003.27$ $32\%$ $63734.97$ $64431.54$ $39\%$
@@ 6.4 Effect of Increasing Penalty Weight
In this experiment, we initialized the penalty weight at $100$, with a growth rate of $1.05$. Table 5 compares the results under this approach. In general, we observe that the method successfully improves the feasibility at the cost of objective value. It reflects the trade-off between enhancing stricter constraint satisfaction and achieving better solutions. However, for the 20000×5 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions.
@@ 6.5 Effect of Training Sample Size
The large number of infeasible solutions observed in the 20000×5 Rosenbrock problem can primarily be attributed to significant overfitting within the model. Given that we have prior knowledge of the parameter distribution and our self-supervised learning approach does not rely on optimal solution labels, we can easily scale up the sample size to effectively mitigate overfitting.
In our experiments with the 20000×5 Rosenbrock problem, we observed significant overfitting due to the large size of the neural network. Given that we have prior knowledge of the parameter distribution and our self-supervised learning approach does not rely on optimal solution labels, we can easily scale up the sample size to effectively mitigate overfitting.
To evaluate the impact of sample size on model performance, we trained our model for 20000×5 Rosenbrock on datasets of three different sizes: 800, 8,000, and 80,000 instances. We adjusted the training epochs to 2000, 200, and 20 (with early stopping), respectively, to maintain a similar number of parameter updates across different dataset sizes. All other hyperparameters were kept consistent with previous experiments to isolate the effect of sample size.
Figure 3: Performance of RC and LT Methods on the 20000×5 Rosenbrock Problem with Varying Training Sample Sizes (800, 8000, 80000).
(a) Infeasible Ratio of RC and LT Methods on Test Set across Different Training Sample Sizes.
(b) Objective Value for Each Instance of RC and LT Methods on Test Set across Different Training Sample Sizes.
As shown in Figure 3, increasing the sample size yields significant improvements in both objective values and feasibility. With 80,000 samples for training, the infeasibility ratio was reduced to 1% on the test set, demonstrating that larger sample sizes help the model generalize better to unseen instances, particularly in extremely high-dimensional settings. These results underscore the importance of adequate sample size when training models. This strategy is especially practical with our approach due to its self-supervised nature, which bypasses the need for labeled optimal solutions.
@@ 7 Conclusion
We have introduced a new learning-based heuristic method for MINLP. Our approach includes two novel correction layers—rounding classification and learnable threshold—that enable neural networks to generate high-quality integer solutions while preserving gradient information for training through backpropagation. These layers allow us to tackle optimization tasks with discrete variables and non-linear constraints in a way that is scalable and computationally efficient. As a self-supervised approach, our method does not require collecting optimal solutions as labels, significantly reducing the time and effort typically needed for data collection. Additionally, we proposed a strategy of progressively increasing penalties to further enhance the feasibility of the predicted solutions.
@@
Our method allows for solving large-scale parametric MINLPs heuristically, leading to significantly better solutions and computational efficiency; however, our methods do not guarantee feasibility. Future work could focus on furthering the feasibility guarantees through alternative constraint handling or solution post-processing. For specific problem classes, it may be possible to relax only a subset of the constraints into the learning loss function while optimizing over the remaining constraints directly using existing differentiable optimization layers, e.g., from Agrawal et al. (2019). Another potential extension of our work is to redesign the neural network architecture to accommodate a varying number of instance parameters and decision variables. This can be achieved by the use of set-based, permutation-equivariant architectures including but not limited to graph neural networks (Cappart et al., 2023; Dumouchelle et al., 2024).
@@ Appendix A Gumbel-Sigmoid Trick
Appendix B Neural Network Structure and Hyperparameters
@@
The constraint penalty weight $\lambda$ was set to $100$. All networks were trained using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $64$ over $200$ epochs. Early stopping was applied based on validation performance to ensure convergence without overfitting.
Appendix C Details for Ablation Study
C.1 Rounding after Learning
Rounding after Learning serves as an ablation study, where direct rounding is applied to the relaxed solution generated by the learning-based model $\pi_{\Theta_{1}}(\cdot)$. In this case, the loss function minimizes the objective based on the continuous relaxed solution $\bar{\bf x}\in\mathbb{R}^{n_{r}+n_{z}}$. However, when direct rounding is applied, the resulting integer solution can cause significant deviations in the objective value and may fail to satisfy feasibility constraints. This highlights the importance of end-to-end learning, where the model receives updates directly from the ultimate loss function.
C.2 Rounding with Straight-through Estimator
As another part of our ablation study, we then implement the STE to handle the non-differentiable rounding operations: 1) In this approach, the relaxed solution $\bar{\bf x}^{i}_{z}$ for integer variables is rounded down, and the STE technique enables effective gradients to pass through the rounding step during backpropagation. 2) To determine the rounding direction ${\bf b}$, we employ the Gumbel-Sigmoid function to provide approximation gradients for backward passes and output a binarized result during the forward passes.

### v4

@@ 6.4 Multi-Dimensional Mixed-Integer Rosenbrock Problem
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Appendix E. Due to the structure of the constraints, it is straightforward to always construct a trivial feasible integer solution, ensuring that the problem remains feasible across instances.
@@
We conducted experiments on mixed-integer Rosenbrock problems with the number of decision variables ranging from 2 to 20,000; the number of constraints was fixed at 5. The results in Table 4 show that RC and LT exhibit strong performance, even outperforming EX in smaller cases. However, as the problem size increases to 10,000 variables, a noticeable decline in feasibility is observed for both RC and LT, while solver-based methods such as EX, N1, and RR fail to produce any solutions. As seen in previous experiments, RR, which relies on rounding relaxations, continues to suffer from significant infeasibility issues.
@@
Table 4: Results for the Mixed-Integer Rosenbrock Problem. The number of decision variables varies from 2 to 20,000, while the number of constraints is 5. See the caption of Table 2 for details. The “—” symbol indicates that no solution is found for any instance within 60 sec time limits. “Unsolved” denotes the percentage of instances that could not be solved within the given time limit.
Method Metric 2×4 20×4 200×4 2000×4 20000×4
RC Obj Mean / Median $23.27/21.48$ $59.39/48.86$ $503.51/461.71$ $5938.37/5792.52$ $66883.20/66797.11$
% Infeasible $3\%$ $0\%$ $1\%$ $1\%$ $24\%$
Time (Sec) $0.0019$ $0.0019$ $0.0021$ $0.0033$ $0.0116$
LT Obj Mean / Median $23.18/20.80$ $62.51/63.40$ $622.78/626.04$ $5611.78/5557.82$ $47622.18/34518.87$
% Infeasible $2\%$ $0\%$ $0\%$ $3\%$ $34\%$
Time (Sec) $0.0019$ $0.0020$ $0.0026$ $0.0030$ $0.0127$
EX Obj Mean / Median $19.62/18.20$ $65.50/59.16$ $911.909.93e5/$ 911.90 $9262.092.50e11/$ 9262.09
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $8\%$ $100\%$
Time (Sec) $3.5630$ $60.2459$ $60.1416$ $60.3500$
RR Obj Mean / Median $22.24/22.19$ $51.171.20e4/$ 51.17 $501.901.43e4/$ 501.90 $85346.617.02e8/$ 85346.61
% Infeasible $45\%$ $41\%$ $18\%$ $1\%$
% Unsolved $0\%$ $0\%$ $42\%$ $92\%$ $100\%$
Time (Sec) $0.1877$ $0.5635$ $1.4163$ $8.4646$
N1 Obj Mean / Median $40.37/27.93$ $87.83/77.34$ $957.423.72e8/$ 957.42 $9325.378.31e12/$ 9325.37
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
% Unsolved $0\%$ $0\%$ $0\%$ $6\%$ $100\%$
Time (Sec) $0.0313$ $0.0829$ $0.2241$ $12.4214$
6.5 Effect of Penalty Weight
This section investigates the impact of the penalty weight, a critical hyperparameter, on the performance of the optimization methods. Experiments were conducted on three representative problems: a 1000×1000 convex quadratic problem, a 1000×1000 simple non-convex problem, and a 20000×4 Rosenbrock problem. For each problem, we evaluated the RC and LT methods under penalty weights of $1$, $5$, $10$, $50$, $100$, $500$, and $1000$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Illustration of the objective value (Top) and proportion of infeasible solutions (Bottom) on the test set. As the penalty weight increases, the fraction of infeasible solutions decreases while the objective value generally deteriorates, as expected.
@@ 6.5 Effect of Penalty Weight
Figure 3 reveals an inherent trade-off between achieving a higher proportion of feasible solutions and maintaining lower objective values. While increasing the penalty weight improves the feasibility rate, it often results in worse objective values. However, for the 20000×4 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions. This limitation is addressed in Section 6.6.
@@ 6.6 Effect of Training Sample Size
The large number of infeasible solutions observed in the 20000×4 Rosenbrock problem can primarily be attributed to significant overfitting within the model. Given that we have prior knowledge of the parameter distribution and our self-supervised learning approach does not rely on optimal solution labels, we can easily scale up the sample size to effectively mitigate overfitting.
To assess the impact of sample size, we trained the model on datasets of 800, 8,000, and 80,000 instances, adjusting training epochs to 2000, 200, and 20 (with early stopping) to ensure comparable iterations. All other hyperparameters remained consistent to isolate the effect of sample size.
Figure 4: Illustration of the objective value (Left) and proportion of infeasible solutions (Right) of 20000×4 Rosenbrock problem on the test set. As the training sample size increases, the fraction of infeasible solutions decreases while the objective value generally deteriorates, as expected.
As shown in Figure 4, increasing the sample size yields significant improvements in both objective values and feasibility. With 80,000 samples for training, the infeasibility ratio was reduced to 5% on the test set, demonstrating that larger sample sizes help the model generalize better to unseen instances, particularly in extremely high-dimensional settings. These results underscore the importance of adequate sample size when training models. This strategy is especially practical with our approach due to its self-supervised nature, which bypasses the need for labeled optimal solutions.
@@ 7 Conclusion
We have introduced a new learning-based heuristic method for MINLP. Our approach includes two novel correction layers—rounding classification and learnable threshold—that enable neural networks to generate high-quality integer solutions while preserving gradient information for training through backpropagation. These layers allow us to tackle optimization tasks with discrete variables and non-linear constraints in a way that is scalable and computationally efficient. As a self-supervised approach, our method does not require collecting optimal solutions as labels, significantly reducing the time and effort typically needed for data collection.
@@
Our method allows for solving large-scale parametric MINLPs heuristically, leading to significantly better solutions and computational efficiency; however, our methods do not guarantee feasibility. Future work could focus on furthering the feasibility guarantees through alternative constraint handling or solution post-processing. For specific problem classes, it may be possible to relax only a subset of the constraints into the learning loss function while optimizing over the remaining constraints directly using existing differentiable optimization layers, e.g., from Agrawal et al. (2019). Another potential extension of our work is to redesign the neural network architecture to accommodate a varying number of instance parameters and decision variables. This can be achieved by the use of set-based, permutation-equivariant architectures including but not limited to graph neural networks (Cappart et al., 2023; Dumouchelle et al., 2024; Chen et al., 2022b; 2024).
@@ Appendix A Gumbel-Sigmoid Trick
Appendix B Details of Correction Layers
The following subsections describe two distinct approaches for designing the correction layer $\varphi_{\Theta_{2}}$; the same network $\pi_{\Theta_{1}}$ is used in both approaches.
Rounding Classification.
Line 6 of Algorithm 1 is the key step in the rounding classification (RC) approach. For the integer variables, RC applies a stochastic soft-rounding to the output ${\bf h}^{i}_{z}$ of the neural network $\delta_{\Theta_{2}}(\bar{\bf x}^{i},{\bm{\xi}}^{i})$, yielding ${\bf b}^{i}\in\{0,1\}^{n_{z}}$. An entry of ${\bf b}^{i}$ determines whether the continuously relaxed value $\bar{\bf x}^{i}_{z}$ of the corresponding variable is rounded down or up. In the backward pass, STE is used in line 4 for the rounding down operation. In line 6, the derivative of the Sigmoid function is used.
Learnable Threshold.
The key steps of the learnable threshold (LT) approach are described in lines 8 and 9 of Algorithm 1. Rather than use Gumbel-Sigmoid for the rounding as in RC, LT learns a vector of per-variable rounding thresholds, ${\bf v}^{i}\in[0,1]^{n_{z}}$, that the Sigmoid generates in line 8 of Algorithm 1. A variable is rounded up if the fractional part of its relaxed value, $(\bar{{\bf x}}^{i}_{z}-\hat{{\bf x}}^{i}_{z})$, exceeds the threshold. The indicator function $\mathbb{I}(\cdot)$ in line 9 produces a binary output in the forward pass. In the backward pass, the gradient is approximated by that of the Sigmoid function with a slope:
Here, the slope is set to 10 to sharpen the Sigmoid function.
Appendix C Example Illustration
Figure 5: Example of the relaxed solutions $\bar{x},\bar{y}$ and the rounding solutions $\hat{x},\hat{y}$ across different epochs of training for the same sample instance using the Rounding Classification approach.
$\min_{x\in\mathbb{R},y\in\mathbb{Z}}$ $\quad(a-x)^{2}+50(y-x^{2})^{2}$
In this formulation, $x$ is a continuous decision variable, and $y$ is an integer decision variable, subject to linear constraints. The instances have parameters $a$ and $b$, which represent the input features to the neural network; for the instance illustrated in Figure 5, these are set to $3.83$ and $6.04$, respectively.
Appendix D Neural Network Structure and Hyperparameters
@@
The constraint penalty weight $\lambda$ was set to $200$ for convex quadratic problems, $100$ for simple non-convex problems and Rosenbrock problems. All networks were trained using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $64$ over $200$ epochs. Early stopping was applied based on validation performance to ensure convergence without overfitting.
Appendix E MINLP Problem Setup and Parameter Sampling
Convex Quadratic Problems
For the convex quadratic problems, $\bm{Q}$ is a diagonal matrix with entries drawn from a uniform distribution over $[0,0.01]$ to guarantee the convexity of the problem. The $\bm{p}$ was generated with entries drawn from a uniform distribution over $[0,0.1]$. The matrix $\bm{A}$ was drawn from a normal distribution with a standard deviation of $0.1$. The parameter $\bm{b}\in\mathbb{R}^{m}$, the right-hand side vector of the inequality constraints, is sampled from a uniform distribution over the range $[-1,1]$.
Simple Non-convex Problems
For the simple non-convex problems, the setup remains consistent with that of the convex quadratic problems. However, an additional parameter $\bm{d}\in\mathbb{R}^{m}$ is introduced, with each element independently sampled from a uniform distribution over $[-0.5,0.5]$. The parameter $\bm{d}$ modifies the constraint matrix $\bm{A}$ by adding $\bm{d}$ to its first column and subtracting $\bm{d}$ from its second column.
Ronsenbrock Problems.
For the Ronsenbrock Problems, the vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance and generated from a standard normal distribution. The parameter $b$ is uniformly distributed over $[1,8]$ for each instance, and the parameter $\mathbf{a}\in\mathbb{R}^{n}$ represents a vector where elements drawn independently from a uniform distribution over $[0.5,4.5]$. The parameters $b$ and $\mathbf{a}$ influence the shape of the feasible region and the landscape of the objective function, serving as input features to the neural network.
Appendix F Ablation Study
Overview.
To better understand the contribution of the correction layers $\varphi_{\Theta_{2}}$, we include two ablation baselines in our experiments:
  • Rounding after Learning (RL): This baseline trains only the first neural network $\pi_{\Theta_{1}}$, which predicts relaxed solutions. Rounding to the nearest integer is applied post-training, meaning that the rounding step does not participate in the training process. This isolates the effect of excluding the corrective adjustments provided by $\varphi_{\Theta_{2}}$. This direct rounding can lead to significant deviations in the objective value and feasibility violations, underscoring the importance of end-to-end learning where updates are guided by the ultimate loss function.
  • Rounding with STE (RS): In the Algorithm 2, continuous values predicted by $\pi_{\Theta_{1}}$ are rounded during training using the Straight-Through Estimator (STE), allowing gradients to pass through the rounding operator. While this mechanism applies a correction to produce integer values by rounding to the nearest integer, it is not learnable and does not adjust the rounding based on the parameter or the relaxation output. Thus, the correction is fixed and solely determined by the nearest-integer rounding, without leveraging additional learning for refinement.